OpenStudy (anonymous):

The sum of the series 1 + (1 + 2 + 4) + (4 + 6 + 9) + (9 + 12 + 16) + ....... + (361 + 380 + 400) is ?

OpenStudy (anonymous):

Use calculator and start adding.. ha ha hah a..

OpenStudy (experimentx):

find the pattern (1 + 2 + 4) + (4 + 6 + 9) + (9 + 12 + 16) + 1 2 3

OpenStudy (anonymous):

Does this Have a pattern?? i am unable

OpenStudy (anonymous):

Of course this is a pattern...

OpenStudy (anonymous):

1 + (7) + (19) + (37) 6 12 18 24 30 like that..

OpenStudy (anonymous):

Every times the number is increasing in multiples of 6..

OpenStudy (anonymous):

@Yahoo! water in eyes got it exactly right. . . its a.p with d=6

OpenStudy (experimentx):

what could be the next term in the series?

OpenStudy (anonymous):

37 + 24 = 61

OpenStudy (experimentx):

(16 + 20 + 25) + (25 + 30 + 36) + ...

OpenStudy (experimentx):

of you look at it ... this is just sum of 3 times n^2 + normal arithmetic series

OpenStudy (anonymous):

61 + 30 = 91

OpenStudy (experimentx):

and you know n^2 = n(n+1)(2n+1)/3 ... just do it that way.

OpenStudy (anonymous):

by 6 i guess..

OpenStudy (anonymous):

yes /6

OpenStudy (experimentx):

and by the way 19^2 = 361 ... can you do it?

OpenStudy (anonymous):

@experimentX i did nt understand

OpenStudy (anonymous):

we figured out the pattern next??

OpenStudy (anonymous):

@ yahoo yes we have in a.p. a=1 by d=6 water in eyes got it perfect

OpenStudy (experimentx):